# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 3. Transposable spaces, trace and Frobenius forms, and a family of tractable metrics
Note for $\mathtt{g}_{1},\mathtt{g}_{2}\in\mathfrak{g}$, $\langle\mathtt{g}_{1},\mathtt{g}_{2}\rangle_{\mathcal{F}}=\langle\mathtt{g}_{1}^{\mathsf{T}},\mathtt{g}_{2}\rangle_{\mathcal{K}}$, hence, if $\langle\mathtt{g}_{1},\mathtt{g}_{2}\rangle_{\mathcal{K}}=0$ for all $\mathtt{g}_{2}\in\mathfrak{g}$, then $\langle\mathtt{g}_{1}^{\mathsf{T}},\mathtt{g}_{2}\rangle_{\mathcal{F}}=0$ for all $\mathtt{g}_{2}$, therefore $\mathtt{g}_{1}^{\mathsf{T}}=0$, or $\mathtt{g}_{1}=0$. Thus, the trace form is nondegenerate on $\mathfrak{g}$.
The next statement is a consequence of $\mathfrak{v}=\mathfrak{v}^{\mathsf{T}}$, which implies $\mathfrak{v}_{{}_{\perp}}=\{\mathtt{h}\in\mathfrak{g}|\;\operatorname{Tr}\mathtt{h}\mathfrak{v}^{\mathsf{T}}=\{0\}\}$, or $\mathfrak{v}_{{}_{\perp}}$ is also the orthogonal complement under the Frobenius inner product. Using $\operatorname{Tr}a^{\mathsf{T}}b=\operatorname{Tr}ab^{\mathsf{T}}$ for two matrices $a,b$, we get $\mathfrak{v}_{{}_{\perp}}$ is transposable. From here, the decomposition $\mathtt{g}=\mathtt{g}_{\mathfrak{v}}+\mathtt{g}_{\mathfrak{v}_{{}_{\perp}}}$ under the trace form is also the decomposition under the Frobenius form. Thus, the projection under one form is the projection under the other. Transposing the decomposition to $\mathtt{g}^{\mathsf{T}}=\mathtt{g}_{\mathfrak{v}}^{\mathsf{T}}+\mathtt{g}_{\mathfrak{v}_{{}_{\perp}}}^{\mathsf{T}}$ then use the uniqueness of this decompostion, we get $(\mathtt{g}^{\mathsf{T}})_{\mathfrak{v}}=(\mathtt{g}_{\mathfrak{v}})^{\mathsf{T}}$.
To show $[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]\subset\mathfrak{a}_{{}_{\perp}}$, let $\mathtt{c}\in\mathfrak{a}$, using eq. 10, we compute $\operatorname{Tr}\mathtt{c}[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]=\operatorname{Tr}\mathfrak{a}_{{}_{\perp}}[\mathtt{c},\mathfrak{a}]=0$ as $[\mathtt{c},\mathfrak{a}]\subset\mathfrak{a}$ since $\mathfrak{a}$ is a subalgebra. For $\mathtt{a}\in\mathfrak{a},\mathtt{b}\in\mathfrak{a}_{{}_{\perp}}$, $[\mathtt{a},\mathtt{b}]^{\mathsf{T}}=-[\mathtt{a}^{\mathsf{T}},\mathtt{b}^{\mathsf{T}}]\subset[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]$, proving $\mathfrak{a}_{{}_{\Join}}^{\mathsf{T}}=\mathfrak{a}_{{}_{\Join}}$. We have eq. 15 implies eq. 14 because $\operatorname{Tr}\mathtt{c}[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]=\operatorname{Tr}\mathfrak{a}_{{}_{\perp}}[\mathtt{c},\mathfrak{a}]$. For the reverse, since $\mathtt{c}\in\mathfrak{a}_{{}_{\perp}}$, $[\mathtt{c},\mathfrak{a}]\in\mathfrak{a}_{{}_{\perp}}$ from eq. 13. As $\mathfrak{a}_{{}_{\perp}}$ is transposable, the trace form is nondegenerate on $\mathfrak{a}_{{}_{\perp}}$, and this shows $[\mathtt{c},\mathfrak{a}]=0$.
Finally, using eq. 15 and $(\mathfrak{a}_{{}_{\top}})_{{}_{\perp}}=\mathfrak{a}\oplus\mathfrak{a}_{{}_{\Join}}$, we get $\mathfrak{a}\mapsto\mathfrak{a}_{{}_{\top}}$ is involutive. ∎
From now on, when $\mathfrak{v}$ is transposable, we will use the notation $\mathtt{p}_{\mathfrak{v}}$ to denote the projection under either the Frobenius or trace form, and write $w_{\mathfrak{v}}$ for $\mathtt{p}_{\mathfrak{v}}(w)$ for $w\in\mathcal{E}$. We denote by $\mathfrak{v}_{{}_{\perp}}$ the orthogonal complement from either the trace or Frobenius form.
@@
Let $\mathrm{G}\subset\mathcal{E}:=\mathbb{R}^{n\times n}$ be a connected matrix Lie group with Lie algebra $\mathfrak{g}$.
@@
Let $\mathrm{G}\subset\mathbb{R}^{n\times n}$ be a connected matrix Lie group with a transposable Lie subalgebra $\mathfrak{g}$, and $\mathfrak{a}\subset\mathfrak{g}$ is a transposable Lie subalgebra. Let $\vec{\beta}=(\beta_{0},\beta_{1})\in\mathbb{R}^{2}$ where $\beta_{0}\beta_{1}\neq 0$, and set $\beta=\frac{\beta_{1}}{\beta_{0}}$. Define the symmetric bilinear form on $\mathfrak{g}$
(17) $\langle\mathtt{g},\mathtt{h}\rangle_{\vec{\beta}}=\beta_{0}(\operatorname{Tr}(\mathtt{h}\mathtt{g})-\operatorname{Tr}(\mathtt{h}_{\mathfrak{a}}\mathtt{g}_{\mathfrak{a}}))-\beta_{1}\operatorname{Tr}(\mathtt{h}_{\mathfrak{a}}\mathtt{g}_{\mathfrak{a}}).$
The form is nondegenerate, which defines a left-invariant pseudo-Riemannian metric on the Lie group $\mathrm{G}$ by eq. 9. For two vector fields $\mathtt{Y},\mathtt{Z}$, the Levi-Civita connection of this metric is given by $\nabla_{\mathtt{Y}}\mathtt{Z}=\mathrm{D}_{\mathtt{Y}}\mathtt{Z}+\Gamma(\mathtt{Y},\mathtt{Z})$ where the Christoffel function $\Gamma$ for two tangent vectors $\xi,\eta\in T_{X}\mathrm{G}$ is given by
(18) $\Gamma(\xi,\eta)=-\frac{1}{2}(\xi X^{-1}\eta+\eta X^{-1}\xi)+\frac{1+\beta}{2}X\left([(X^{-1}\xi)_{\mathfrak{a}},X^{-1}\eta]+[(X^{-1}\eta)_{\mathfrak{a}},X^{-1}\xi]\right).$
Further, the last term above is in $X\mathfrak{a}_{{}_{\Join}}$, as for $\mathtt{a},\mathtt{b}\in\mathfrak{g}$, we have
(19) $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}]+[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\in\mathfrak{a}_{{}_{\Join}}.$
The geodesic $\gamma(t)$ starting at $\gamma(0)=X\in\mathrm{G}$ with initial velocity $\dot{\gamma}(0)=\xi$ is given by
(20) $\gamma(t)=X\operatorname{expm}(t\left(-(1+\beta)(X^{-1}\xi)_{\mathfrak{a}}+X^{-1}\xi\right))\operatorname{expm}(t(1+\beta)(X^{-1}\xi)_{\mathfrak{a}}).$
Set $\mathtt{a}=X^{-1}\xi$, let $\eta\in T_{X}\mathcal{M}$ be another tangent vector, then the parallel transport of $\eta$ along the geodesics $\gamma$ to $X_{t}=\gamma(t)$ with initial point and velocity $(X,\xi)$ is
@@
(21) $\begin{gathered}\mathtt{T}(\gamma)_{0}^{t}\eta=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))\operatorname{expv}(t\mathrm{P}_{\mathtt{a}},X^{-1}\eta)\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})\\ \text{ where for }\mathtt{b}\in\mathfrak{g},\mathrm{P}_{\mathtt{a}}:\mathtt{b}\mapsto\frac{1}{2}\left([\mathtt{b},\mathtt{a}]+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}])\right).\end{gathered}$
In the Frobenius inner product, the adjoint of the operator $\mathrm{P}_{\mathtt{a}}$ is given by
(22) $\mathrm{P}_{\mathtt{a}}^{\mathsf{T}}\mathtt{b}=\frac{1}{2}\left([\mathtt{b},\mathtt{a}^{\mathsf{T}}]+(1+\beta)([\mathtt{a}_{\mathfrak{a}}^{\mathsf{T}},\mathtt{b}]-[\mathtt{b},\mathtt{a}^{\mathsf{T}}]_{\mathfrak{a}})\right).$
The operator $\mathrm{P}_{\mathtt{a}}$ is antisymmetric in the bilinear pairing eq. 17. For $\mathtt{a},\mathtt{b},\mathtt{c}\in\mathfrak{g}$

### v2

@@ 3. The Trace form as a bi-invariant symmetric bilinear form and transposable subalgebras
Note for $\mathtt{A}_{1},\mathtt{A}_{2}\in\mathfrak{g}$, $\langle\mathtt{A}_{1},\mathtt{A}_{2}\rangle_{\mathscr{F}}=\langle\mathtt{A}_{1}^{\mathsf{T}},\mathtt{A}_{2}\rangle_{\mathscr{K}}$, hence, if $\langle\mathtt{A}_{1},\mathtt{A}_{2}\rangle_{\mathscr{K}}=0$ for all $\mathtt{A}_{2}\in\mathfrak{g}$, then $\operatorname{Tr}\mathtt{A}_{1}^{\mathsf{T}}\mathtt{A}_{2}=0$ for all $\mathtt{A}_{2}$, therefore $\mathtt{A}_{1}=0$. Thus, the trace form is nondegenerate on $\mathfrak{g}$.
The next statement is a consequence of $\mathfrak{v}=\mathfrak{v}^{\mathsf{T}}$, which implies $\mathfrak{v}_{{}_{\perp}}=\{\mathtt{A}\in\mathfrak{g}|\;\operatorname{Tr}\mathtt{A}\mathfrak{v}^{\mathsf{T}}=\{0\}\}$, or $\mathfrak{v}_{{}_{\perp}}$ is also the orthogonal complement under the Frobenius inner product. Using $\operatorname{Tr}\mathtt{A}^{\mathsf{T}}\mathtt{B}=\operatorname{Tr}\mathtt{A}\mathtt{B}^{\mathsf{T}}$ for two matrices $\mathtt{A},\mathtt{B}$, we get $\mathfrak{v}_{{}_{\perp}}$ is transposable. From here, the decomposition $\mathtt{A}=\mathcal{P}_{\mathfrak{v}}\mathtt{A}+\mathcal{P}_{\mathfrak{v}_{{}_{\perp}}}\mathtt{A}$ under the trace form is also the decomposition under the Frobenius form. Thus, the projection under one form is the projection under the other. Transposing the decomposition to $\mathtt{A}^{\mathsf{T}}=(\mathcal{P}_{\mathfrak{v}}\mathtt{A})^{\mathsf{T}}+(\mathcal{P}_{\mathfrak{v}_{{}_{\perp}}}\mathtt{A})^{\mathsf{T}}$ then use the uniqueness of this decomposition, we get $\mathcal{P}_{\mathfrak{v}}(\mathtt{A}^{\mathsf{T}})=(\mathcal{P}_{\mathfrak{v}}\mathtt{A})^{\mathsf{T}}$. ∎
@@
4. Pseudo-Riemannian deformation metric
The main result of this section is if $\mathcal{G}$ has a non-degenerate bi-invariant pseudo-Riemannian metric $\langle,\rangle$, and $\mathfrak{a}\subset\mathfrak{g}$ is a nondegenerate subalgebra, if we replace the bilinear form on $\mathfrak{a}$ with minus of another bi-invariant form $\langle,\rangle_{\mathcal{B}}$ in $\mathfrak{a}$ while scaling the form on the complement $\mathfrak{a}_{{}_{\perp}}$ by a scalar $\beta_{0}\neq 0$, then the resulting left-invariant metric on $\mathcal{G}$ will be naturally reductive (but may no longer be bi-invariant) and we still have closed-form geodesics and tractable parallel transport. Relative to the original form, the form $\langle,\rangle_{\mathcal{B}}$ on $\mathfrak{a}$ is defined by a symmetric (with respect to $\langle,\rangle$) linear operator $\mathcal{B}$ on $\mathfrak{a}$ only ($\mathcal{B}\mathtt{A}$ only makes sense if $\mathtt{A}\in\mathfrak{a}$), called an intertwining operator. Thus, the new metric will be given by
(13) $\langle\mathtt{A},\mathtt{A}\rangle_{{}_{(\beta_{0},\mathcal{B})}}:=\beta_{0}\langle\mathcal{P}_{\mathfrak{a}_{{}_{\perp}}}(\mathtt{A}),\mathcal{P}_{\mathfrak{a}_{{}_{\perp}}}(\mathtt{A})\rangle-\langle\mathcal{P}_{\mathfrak{a}}(\mathtt{A}),\mathcal{B}\mathcal{P}_{\mathfrak{a}}(\mathtt{A})\rangle.$
We first discuss the condition for $\langle,\rangle_{\mathcal{B}}$ to be bi-invariant.
4.1. Intertwining operator
We note that the most important examples of intertwining operators we use are simply multiplying by scalars, so in the procedure above, we just scale the metrics on $\mathfrak{a}$ and $\mathfrak{a}_{{}_{\perp}}$ by possibly different scalars. However, the general result requires roughly the same effort to prove, and potentially could have future applications.
If a Lie algebra $\mathfrak{a}$ has a nondegenerate bi-invariant symmetric form $\langle,\rangle$, the form is not unique, as any scalar multiple of $\langle,\rangle$ is also bi-invariant. If $\mathfrak{a}$ is a direct sum of two mutually commuting subalgebras $\mathfrak{a}=\mathfrak{a}_{1}\oplus\mathfrak{a}_{2}$, then the form $(\mathtt{A},\mathtt{B})\mapsto\beta_{1}\langle\mathtt{A}_{1},\mathtt{B}_{1}\rangle+\beta_{2}\langle\mathtt{A}_{2},\mathtt{B}_{2}\rangle,\mathtt{A},\mathtt{B}\in\mathfrak{a}$ is also bi-invariant, where $\mathtt{A}=\mathtt{A}_{1}+\mathtt{A}_{2},\mathtt{B}=\mathtt{B}_{1}+\mathtt{B}_{2}$ is the decomposition of $\mathtt{A}$ and $\mathtt{B}$ to the $\mathfrak{a}_{1}$ and $\mathfrak{a}_{2}$ components. Here, we apply different scalar factors on different components. We note if $\mathfrak{a}$ is commutative, $\mathtt{A}_{1},\mathtt{A}_{2}]=0$ for all $\mathtt{A}_{1},\mathtt{A}_{2}\in\mathfrak{a}$ then any form is bi-invariant. These are the often used constructions of bi-invariant forms, see also [27, Lemma 7.5,7.6], [26, 28].
We now give a condition for an operator to preserve the bi-invariant property. It satisfies the same condition as intertwining operators (morphism of the adjoint representation [15, section II.6.1]) in representation theory. Thinking of them as generalized scalars would probably be helpful. Most material presented here is likely known, but we could not locate the references.
Note that if $\langle,\rangle$ is non-degenerate, any other symmetric form could be written as $\langle\mathtt{A},\mathtt{C}\rangle_{\mathcal{B}}:=\langle\mathtt{A},\mathcal{B}\mathtt{C}\rangle$ ($\mathtt{A},\mathtt{C}\in\mathfrak{a}$) for a symmetric operator $\mathcal{B}$ (that means $\langle\mathcal{B}\mathtt{A},\mathtt{C}\rangle=\langle\mathtt{A},\mathcal{B}\mathtt{C}\rangle$). Also note that $\mathcal{B}$ is invertible if and only if $\langle,\rangle_{\mathcal{B}}$ is non-degenerate.
Definition 2.
Let $\mathfrak{a}$ be a Lie algebra with a nondegenerate bi-invariant form $\langle,\rangle$, then an intertwining operator $\mathcal{B}$ is a symmetric operator on $\mathfrak{a}$ such that the form
(14) $(\mathtt{A}_{1},\mathtt{A}_{2})\mapsto\langle\mathtt{A}_{1},\mathtt{A}_{2}\rangle_{\mathcal{B}}:=\langle\mathtt{A}_{1},\mathcal{B}\mathtt{A}_{2}\rangle$
is also bi-invariant.
We have the following characterization of intertwining operators
Proposition 2.
A symmetric operator $\mathcal{B}$ is an intertwining operator if and only if $\mathcal{B}$ satisfies $\mathcal{B}\operatorname{ad}_{\mathtt{A}}=\operatorname{ad}_{\mathtt{A}}\mathcal{B}$ for all $\mathtt{A}\in\mathfrak{a}$. Equivalently, it is an intertwining operator if either of the following holds
(15) $\mathcal{B}[\mathtt{A}_{1},\mathtt{A}_{2}]=[\mathtt{A}_{1},\mathcal{B}\mathtt{A}_{2}],$
(16) $\mathcal{B}[\mathtt{A}_{1},\mathtt{A}_{2}]=[\mathcal{B}\mathtt{A}_{1},\mathtt{A}_{2}].$
Proof.
For $\langle,\rangle_{\mathcal{B}}$ to be bi-invariant, we need $\langle\mathtt{A}_{1},[\mathtt{A}_{2},\mathtt{A}_{3}]\rangle_{\mathcal{B}}=\langle\mathtt{A}_{2},[\mathtt{A}_{3},\mathtt{A}_{1}]\rangle_{\mathcal{B}}$, for $\mathtt{A}_{1},\mathtt{A}_{2},\mathtt{A}_{3}\in\mathfrak{a}$ or
$\langle\mathtt{A}_{1},\mathcal{B}[\mathtt{A}_{2},\mathtt{A}_{3}]\rangle=\langle\mathtt{A}_{2},\mathcal{B}[\mathtt{A}_{3},\mathtt{A}_{1}]\rangle$
but the first term could be written as $\langle\mathcal{B}\mathtt{A}_{1},[\mathtt{A}_{2},\mathtt{A}_{3}]\rangle=\langle\mathtt{A}_{2},[\mathtt{A}_{3},\mathcal{B}\mathtt{A}_{1}]\rangle$. This holds for all $\mathtt{A}_{2}$, and since $\langle,\rangle$ is nondegenerate, $\mathtt{A}_{3},\mathcal{B}\mathtt{A}_{1}]=\mathcal{B}[\mathtt{A}_{3},\mathtt{A}_{1}]$, or $\mathcal{B}$ commutes with $\operatorname{ad}_{\mathtt{A}_{3}}$
(17) $\operatorname{ad}_{\mathtt{A}}\mathcal{B}=\mathcal{B}\operatorname{ad}_{\mathtt{A}}\quad\text{ for }\mathtt{A}\in\mathfrak{a}.$
Finally, note the operator $\mathtt{X}\mapsto[\mathtt{X},\mathtt{A}]$ is $\operatorname{ad}_{\mathtt{A}}$, thus, if one of eq. 15 or eq. 16 holds for all $\mathtt{A}_{1},\mathtt{A}_{2}$ in $\mathfrak{a}$, we have the other. ∎
In the case of $\mathfrak{so}(n)$, or of a class of Lie algebras called simple, $\mathcal{B}$ must be a scalar multiplication. Most intertwining operators in this paper are of this form. The following shows intertwining operators behave like scalars.
Lemma 2.
Let $\operatorname{I}_{\mathfrak{a}}$ be the identity operator on $\mathfrak{a}$, then $\beta\operatorname{I}_{\mathfrak{a}},\beta\in\mathbb{R}$, is an intertwining operator. Sum, product, and inversion (if exists) of intertwining operators are intertwining operators.
For example, if $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are intertwining operators then
$\operatorname{ad}_{\mathtt{A}}\mathcal{B}_{1}\mathcal{B}_{2}=\mathcal{B}_{1}\operatorname{ad}_{\mathtt{A}}\mathcal{B}_{2}=\mathcal{B}_{1}\mathcal{B}_{2}\operatorname{ad}_{\mathtt{A}},$
and $\operatorname{ad}_{\mathtt{A}}\mathcal{B}(\mathcal{B}^{-1}\mathtt{A}_{1})=\mathcal{B}\operatorname{ad}_{\mathtt{A}}(\mathcal{B}^{-1}\mathtt{A}_{1})$ implies $\mathcal{B}^{-1}\operatorname{ad}_{\mathtt{A}}\mathtt{A}_{1}=\operatorname{ad}_{\mathtt{A}}(\mathcal{B}^{-1}\mathtt{A}_{1})$, proving product and inversion producing new intertwining operators.
4.2. Deformation using an intertwining operator on a subalgebra
We will consider matrix groups in the following theorem to simplify the notation, but we believe it also holds for abstract Lie groups, replacing matrix multiplications with left and right multiplication actions.
Proposition 3.
Let $\mathcal{G}$ be a connected matrix Lie group with Lie algebra $\mathfrak{g}$ and a nondegenerate bi-invariant form $\langle,\rangle$. Let $\mathcal{A}$ be a connected Lie subgroup of $\mathcal{G}$ with Lie algebra $\mathfrak{a}\subset\mathfrak{g}$ such that $\langle,\rangle$ restricted to $\mathfrak{a}$ is nondegenerate. Let $\mathcal{B}$ be an invertible intertwining operator on $\mathfrak{a}$. Write $\vec{\beta}$ for the tuple $(\beta_{0},\mathcal{B})$ where $\beta_{0}\neq 0$ is a real number. For $\mathtt{X}_{1},\mathtt{X}_{2}\in\mathfrak{g}$, define another bilinear form on $\mathfrak{g}$
(18) $\langle\mathtt{X}_{1},\mathtt{X}_{2}\rangle_{\vec{\beta}}:=\beta_{0}\langle\mathtt{X}_{1},\mathtt{X}_{2}\rangle-\langle(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle.$
It induces a left-invariant metric on $\mathcal{G}$. Let $\tilde{\mathcal{G}}$ be the component-wise product $\mathcal{G}\times\mathcal{A}$ and $\tilde{\mathcal{H}}=\{(H,H)\in\tilde{\mathcal{G}}|H\in\mathcal{A}\}$. Then the action $(X,A).Y=XYA^{-1}$ makes $\tilde{\mathcal{G}}$ a group of isometries of $\mathcal{G}$ with the metric eq. 18. The isotropy group at $\operatorname{I}_{\mathcal{G}}$ is $\tilde{\mathcal{H}}$. Thus, we can identify $\mathcal{G}$ with $\tilde{\mathcal{G}}$$\tilde{\mathcal{H}}$.
Proof.
First, eq. 18 is equivalent to eq. 13 since $\mathcal{P}_{\mathfrak{a}_{{}_{\perp}}}=\operatorname{I}_{\mathfrak{g}}-\mathcal{P}_{\mathfrak{a}}$, and $\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathtt{X}_{2}\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle$. The form $\langle,\rangle$ is nondegenerate on $\mathfrak{a}_{{}_{\perp}}\subset\mathfrak{g}$ by lemma 1. If $\mathtt{X}_{1}$ is a degenerate element in eq. 13, it pairs to zero with all $\mathtt{X}_{2}\in\mathfrak{g}$. For $\mathtt{X}_{2}$ in $\mathfrak{a}$, this implies $\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1}=0$, while for $\mathtt{X}_{2}\in\mathfrak{a}_{{}_{\perp}}$, we get $\mathcal{P}_{\mathfrak{a}_{{}_{\perp}}}\mathtt{X}_{1}=0$. Together, we have $\mathtt{X}_{1}=0$. Thus, the form eq. 18 is nondegenerate.
By definition, the metric eq. 18 is left invariant. To show for $A\in\mathcal{A}$, $Y\mapsto YA^{-1}$ is an isometry on $\mathcal{G}$, the first term of eq. 18 is bi-invariance. We need to deal with the second term. We have $\operatorname{Ad}_{A}$ commutes with $\mathcal{P}_{\mathfrak{a}}$ since for $\mathtt{X}_{1},\mathtt{X}_{2}\in\mathfrak{g}$,
$\begin{gathered}\langle\mathcal{P}_{\mathfrak{a}}\operatorname{Ad}_{A}\mathtt{X}_{1},\mathtt{X}_{2}\rangle=\langle\operatorname{Ad}_{A}\mathtt{X}_{1},\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle=\langle\mathtt{X}_{1},\operatorname{Ad}_{A^{-1}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\operatorname{Ad}_{A^{-1}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle\\ =\langle\operatorname{Ad}_{A}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2}\rangle=\langle\operatorname{Ad}_{A}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathtt{X}_{2}\rangle.\end{gathered}$
Thus, $\mathcal{P}_{\mathfrak{a}}\operatorname{Ad}_{A}\mathtt{X}_{1}=\operatorname{Ad}_{A}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1}$. Since $\mathcal{A}$ is connected, by eq. 6, $\mathcal{B}$ commutes with $\operatorname{Ad}$. Hence, for $\xi,\eta\in T_{Y}\mathcal{G}$, we have $\xi A^{-1},\eta A^{-1}\in T_{YA^{-1}}\mathcal{G}$ and we need to evaluate
$\begin{gathered}\langle(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}\left((YA^{-1})^{-1}\xi A^{-1}\right),\mathcal{P}_{\mathfrak{a}}\left((YA^{-1})^{-1}\eta A^{-1}\right)\rangle\\ =\langle(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}\left(A(Y^{-1}\xi)A^{-1}\right),\mathcal{P}_{\mathfrak{a}}\left(A(Y^{-1}\eta)A^{-1}\right)\rangle\\ =\langle A(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}(Y^{-1}\xi)A^{-1},A\mathcal{P}_{\mathfrak{a}}(Y^{-1}\eta)A^{-1}\rangle=\langle(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}(Y^{-1}\xi),\mathcal{P}_{\mathfrak{a}}(Y^{-1}\eta)\rangle\end{gathered}$
by $\operatorname{Ad}_{A}$ invariance. Thus, $\tilde{\mathcal{G}}$ acts as an isometry. The isotropy group at $\operatorname{I}_{\mathcal{G}}$ consists of $(X,A)$ such that $XA^{-1}=\operatorname{I}_{\mathcal{G}}$, thus $X=A$, proving the isotropy group is $\tilde{\mathcal{H}}$. Observe for $X\in\mathcal{G}$ and $A\in\mathcal{A}$, the action of $\tilde{\mathcal{G}}$ by $(XA,A)$ sends $\operatorname{I}_{\mathcal{G}}$ to $XAA^{-1}=X$, thus the action is transitive. The inverse image at each $X$ is a copy of $\mathcal{A}$, embedded as $\tilde{\mathcal{H}}$. This identifies $\mathcal{G}$ with $\tilde{\mathcal{G}}$$\tilde{\mathcal{H}}$ as manifolds. ∎
@@ 4.2. Deformation using an intertwining operator on a subalgebra
The metric in proposition 3 is naturally reductive when $\mathcal{G}$ is identified with $\tilde{\mathcal{G}}$$\tilde{\mathcal{H}}$. Specifically, let $\tilde{\mathfrak{g}}=\mathfrak{g}\times\mathfrak{a}$ be the Lie algebra of $\tilde{\mathcal{G}}$ and $\tilde{\mathfrak{h}}$ be the Lie algebra of $\tilde{\mathcal{H}}$. Let $\hat{\mathcal{B}}:=(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}$. Then
(19) $\begin{gathered}\tilde{\mathfrak{h}}=\{(\mathtt{A},\mathtt{A})\in\tilde{\mathfrak{g}}|\mathtt{A}\in\mathfrak{a}\},\\ \tilde{\mathfrak{m}}:=\{(\mathtt{X},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X})|\mathtt{X}\in\mathfrak{g}\}\text{ with }\hat{\mathcal{B}}=(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\end{gathered}$
are complementary subspaces, $\tilde{\mathfrak{g}}=\tilde{\mathfrak{m}}+\tilde{\mathfrak{h}}$ and $\tilde{\mathfrak{m}}\cap\tilde{\mathfrak{h}}=\{0\}$ with the decomposition
@@
(20) $\begin{gathered}(\mathtt{X},\mathtt{A})=\left(\mathtt{X}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}+\beta_{0}^{-1}\mathcal{B}\mathtt{A},-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A})\right)\\ +\left(\mathtt{A}+(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A}),\mathtt{A}+(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A})\right).\end{gathered}$
The space $\tilde{\mathfrak{m}}$ is $\operatorname{Ad}_{\tilde{\mathcal{H}}}$ invariance. The following bilinear form on $\tilde{\mathfrak{m}}$